# Doppler Correspondence: Non-Iterative Scan Matching With Doppler Velocity-Based Correspondence

### v1

@@ VII Conclusion
Overall, the results indicate that Doppler Correspondence does not always guarantee better performance compared with closest point method. Scenarios with high angular velocity may cause the violation of the aforementioned assumption and degradation of odometry performance. However, the proposed method offers distinct advantages, such as its weak dependence on the environment’s geometry, which sets it apart from closest point correspondence. These features enable robust and efficient correspondence estimation, making it a valuable complement to traditional methods in specific applications. Furthermore, its simplicity makes it easy to use and integrate with other methods. While further improvements are needed for broader applicability, the proposed method holds novelty as it leverages Doppler information in correspondence for the first time.

### v2

@@ VII Conclusion
Overall, the results indicate that Doppler Correspondence does not always guarantee better performance compared with the closest point method. Scenarios with high angular velocity may cause the violation of the aforementioned assumption and degradation of odometry performance. However, the proposed method offers distinct advantages, such as its weak dependence on the environment’s geometry, which sets it apart from closest point correspondence. These features enable robust and efficient correspondence estimation, making it a valuable complement to traditional methods in specific applications. Furthermore, its simplicity makes it easy to use and integrate with other methods. While further improvements are needed for broader applicability, the proposed method holds novelty as it leverages Doppler information in correspondence for the first time.
TABLE II: Odometry performance on the semi-dense 4D radar dataset. Each cell reports the absolute pose errors [5] (in meters). Best results are in bold, second-best are underlined.
Method Sequence
loop1_0 loop1_1 loop1_3 loop1_4
ICP point-to-point 61.72 38.71 146.65 3.15
ICP point-to-plane 55.18 43.40 127.52 2.89
Generalized-ICP [17] 64.21 57.28 138.41 3.30
VGICP [9] 50.86 61.86 141.37 4.09
DICP [7] 60.15 68.60 143.24 2.94
Doppler Correspondence (Ours) 54.87 50.04 101.74 7.49
Doppler 4D-ICP (Ours) 63.25 37.74 114.76 2.80
TABLE III: Odometry performance on the dense 4D LiDAR dataset. Each cell reports the absolute pose errors [5] (in meters). Best results are in bold, second-best are underlined. Note that Div. indicates the trajectory of the algorithm diverged.
Method Sequence
Baker-Barry Tunnel
Baker-Barry Tunnel
ICP point-to-point Div. Div.
ICP point-to-plane Div. Div.
Generalized-ICP [17] Div. Div.
VGICP [9] Div. Div.
DICP [7] 65.27 37.12
Doppler Correspondence (Ours) 20.16 47.54
Doppler 4D-ICP (Ours) 14.50 24.06
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grants funded by the MSIT (No. 2023R1A2C2003130) and the Ministry of Education (No. 2020R1A6A1A03040570).